You pour 100 g of boiling water into a 3 kg cast-iron pot at room temperature (20°?). What will the water cool to?
Answer #1

using principle of calorimetry

heat released by water = heat absorbed by iron cast

mw Cw (100 - T) = mi Ci ( T - 20)

0.100 * 4186* ( 100 - T) = 3 * 448 ( T - 20)

31.146 - 0.311 T = T - 20

T = 39

======

Comment before rate in case any doubt, will reply for sure.. goodluck